An anstranaut on the moon picks up a rock and holds it out. At that moment, the rock is at rest and has 5.00 joules of gravitational potential energy. A moment later the rock hit the surface of the moon. What was the kinetic energy of the rock immediately before it hits the surface of the moon?


1
Expert's answer
2019-09-23T09:27:10-0400

Total energy of the rock is the sum of kinetic and potential energies: "E = T + U = const" (this quantity is conserved).

In the first state, the rock is at rest (has zero kinetic energy) and has gravitational potential energy "5 J", and just before it hits the surface of the moon, it has zero potential energy, and some kinetic energy. By the law of the conservation of total energy, "0 + 5 J = T + 0", so just before the rock hits the surface, it has kinetic energy "5 J".